OpenStudy (anonymous):

Factor completely x2 –x–6?

OpenStudy (anonymous):

you need factors of -6 that equal -1, i.e., -3*2=-6 and -3+3=-1, then\[x^2-x-6=(x-3)(x+2)\]

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

The answer is (x+2)(x-3) The reason for this is we need to ask ourselves, what two numbers add to make -1 and multiply together to make -6? The solution -3 and +2. -3 + 2 = -1 -3 * -2 = -6 So we have (x+2)(-3)

OpenStudy (anonymous):

i get sean thanks

OpenStudy (anonymous):

NP.